# AReferences

### v1

@@ Abstract
Identifying the training data samples that most influence a generated image is a critical task in understanding diffusion models, yet existing influence estimation methods are constrained to small-scale or LoRA-tuned models due to computational limitations. As diffusion models scale up, these methods become impractical. To address this challenge, we propose DMin (Diffusion Model influence), a scalable framework for estimating the influence of each training data sample on a given generated image. By leveraging efficient gradient compression and retrieval techniques, DMin reduces storage requirements from 339.39 TB to only 726 MB and retrieves the top-$k$ most influential training samples in under 1 second, all while maintaining performance. Our empirical results demonstrate DMin is both effective in identifying influential training samples and efficient in terms of computational and storage requirements.
Figure 1: Examples of training data samples, shown from high to low influence from left to right, with prompts displayed below each generated image. Samples with higher influence are more similar to the generated image. This examples are from the SD 3 Medium (LoRA) with compressed dimension $v=2^{16}$.
@@ 1 Introduction
Diffusion models have emerged as powerful generative models, capable of producing high-quality images and media across various applications [8, 42, 25, 44]. Despite their impressive performance, the extremely large scale and complexity of the datasets used for training are often sourced broadly from the internet [37, 40, 38, 23]. This vast dataset diversity allows diffusion models to generate an extensive range of content, enhancing their versatility and adaptability across multiple domains [19, 5]. However, it also means that these models may inadvertently generate unexpected or even harmful content, reflecting biases or inaccuracies present in the training data.
This raises an important question: given a generated image, can we estimate the influence of each training data sample for a diffusion model? Such an estimation is crucial for various applications, such as understanding potential biases [17, 26] and improving model transparency by tracing the origins of specific generated outputs [16, 7, 13].
Recently, many studies have explored influence estimation in diffusion models [29, 18, 28, 30, 11]. These methods assign an influence score to each training data sample relative to a generated image, quantifying the extent to which each sample impacts the generation process. For instance, DataInf [18] and K-FAC [28] are influence approximation techniques tailored for diffusion models. However, they are both second-order methods that require the inversion of the Hessian matrix. To approximate this inversion, they must load all the gradients of training data samples across several predefined timesteps. Notably, in the case of the full-precision Stable Diffusion 3 medium model [10], the gradient of entire model requires approximately 8 GB of storage. Collecting gradients for one training sample over 10 timesteps would consume $8\times 10=80$ GB. Scaling this requirement to a training dataset of $10,000$ samples results in a storage demand of around 800 TB – far exceeding the capacity of typical memory or even hard drives. Given that diffusion models are often trained on datasets with millions of samples, this storage demand becomes entirely impractical. Consequently, these methods are limited to LoRA adapters and small diffusion models such as Denoising Diffusion Probabilistic Models (DDPM) [15] or Latent Diffusion Models (LDM) [35], restricting their scalability.
@@
To address these challenges, in this paper, we propose DMin, a scalable influence estimation framework for diffusion models. Unlike existing approaches that are limited to small models or LoRA-tuned models, the proposed DMin scales effectively to larger diffusion models with full parameters. For each data sample, DMin first computes and collects gradients at each timestep, then compresses these gradients from around 40 GB to just 80 KB while maintaining performance. Following this compression, DMin can accurately estimate the influence of each training data sample on a given generated image or retrieve the top-$k$ most influential samples on-the-fly using K-nearest neighbors (KNN) search, enabling further speedup based on the specific task.
@@
  • We introduce DMin, a scalable influence estimation framework for diffusion models, compatible with various architectures, from small diffusion models and LoRA-tuned models to large-scale models like Stable Diffusion.
@@ 2 Influence Estimation for Diffusion Models
Given a test generation $x^{s}$, where $x^{s}$ is generated by a well-trained diffusion model with parameters $\theta^{*}$, the goal of influence estimation is to estimate the influence of each training data sample $x^{i}(1\leq i\leq N)$ on generating $s$, where $N$ is the size of training dataset. Let $z^{i}_{t}$ represents the latent representation of $x^{i}$ on timestep $t$, $z^{s}_{t}$ denotes the latent representation of test generation $s$ on timestep $t$. Considering in the $\alpha$-th training iteration, the model parameter $\theta_{\alpha+1}$ are updated from $\theta_{\alpha}$ by gradient descent on the noise prediction loss for a random timestep $t$ by:
$\theta_{\alpha+1}=\theta_{\alpha}-\eta_{\alpha}\nabla_{\theta_{\alpha}}\mathcal{L}\Bigl{(}f_{\theta_{\alpha}}(z^{i}_{t},t),\epsilon^{i}_{t}\Bigr{)}$ (2)
where $\eta$ denotes the learning rate in the $\alpha$-th iteration. The influence of this training iteration $z^{i}_{t}$ with respect to $z^{s}_{t}$ should be:
@@
To estimate the influence of training data sample $x^{i}$, we can summing up all the training iteration training on $x^{i}$:
@@
where $e$ is the number of epochs, $\bar{\eta}$ is the average learning rate during training, $\epsilon$ is the Gaussian noise for influence estimation, which is the same for all training data samples in estimation. Similarly, for a text-to-image model, the influence of a training data sample $X^{i}=(p^{i},x^{i})$ with respect to test generation $X^{s}=(p^{s},x^{s})$ can be estimated by:
@@ 3.1 Gradient Computation
Since the training dataset remains fixed after training, we can cache the loss gradient of each training data sample, as illustrated in Figure 2(a). For a given training pair $X^{i}=(p^{i},x^{i})$, and a timestep $t$, the training data is processed through the diffusion model in the same way as during training, and a loss is computed between the model-predicted noise and a predefined Gaussian noise $\epsilon$. Back-propagation is then performed to obtain the gradient $g^{i}_{t}$ for the training data pair $X^{i}$ at timestep $t$. Once all gradients $\{g^{i}_{1},g^{i}_{2},\cdots,g^{i}_{T}\}$ for $X^{i}$ at all timesteps are obtained, we apply a compression technique to these gradients and cache the compressed versions for influence estimation. Furthermore, for tasks where only the top-$k$ most influential samples are required, we can construct a KNN index on the compressed gradients to enable efficient querying.
Forward and Backward Passes. In forwarding, following the same process as training, for a training pair $(p^{i},x^{i})$ and a timestep $t$, the prompt $p^{i}$ is passed through the encoder to obtain a prompt embedding, while the image $x^{i}$ is passed through a VAE to obtain a latent representation $z_{0}^{i}$. Gaussian noise is then added to the latent representation to obtain noisy latent presentation. Importantly, as described in Equation 7 and 8, we use the same Gaussian noise $\epsilon$ consistently for all training data samples across all timesteps. To ensure consistency, at the start of the gradient computation process, we sample a single Gaussian noise and use it for all training data samples as well as for the generated image in influence estimation.
@@
Index Construction for KNN. To further enhance the scalability of DMin, we introduce KNN search [45, 39, 41] for tasks requiring only the top-$k$ most influential samples. After gradient compression, as shown in Figure 2(d), we concatenate all the compressed gradients across timesteps to construct a KNN index, enabling efficient querying during influence estimation. This approach is well-suited for extremely large datasets, allowing for the retrieval of the top-$k$ most influential samples on the fly.
@@ 4 Experiments
Models. For conditional text-to-image diffusion model, we use three different models: (1) SD 1.4 with LoRA [^1] CompVis/stable-diffusion-v1-4, (2) SD 3 Medium with LoRA333stabilityai/stable-diffusion-3-medium-diffusers and (3) SD 3 Medium33footnotemark: 3 (Full parameters). For unconditional diffusion model, we conduct experiments on two DDPM models trained on MNIST2221aurent/ddpm-mnist and CIFAR-10444google/ddpm-cifar10-32 respectively. The detailed settings of models are included in Appendix 7.1.
@@
Ours (w/o Comp.) 0.9846 0.9564 0.7574 0.5438 0.7176 0.6176 0.2494 0.1435 0.9821 0.9791 0.9656 0.9639
Ours ($v=2^{12}$, Exact) 0.9846 0.9538 0.7512 0.5233 0.7059 0.6118 0.2388 0.1394 0.9583 0.9425 0.9315 0.9323
Ours ($v=2^{16}$, Exact) 0.9846 0.9564 0.7554 0.5446 0.7294 0.6235 0.2541 0.1459 0.9805 0.9778 0.9666 0.9640
Ours ($v=2^{12}$, KNN) 0.9846 0.9538 0.7513 0.5290 0.6941 0.6176 0.2400 0.1412 0.9604 0.9463 0.9356 0.9364
Ours ($v=2^{16}$, KNN) 0.9846 0.9564 0.7569 0.5492 0.7294 0.6235 0.2506 0.1441 0.9870 0.9840 0.9731 0.9724
@@
Ours (w/o Comp.) 0.9487 0.9436 0.9174 0.8592 0.8706 0.7941 0.4482 0.2718 0.9919 0.9900 0.9828 0.9775
Ours ($v=2^{12}$, Exact) 0.9436 0.9385 0.9097 0.8510 0.8706 0.7706 0.4494 0.2718 0.9919 0.9897 0.9817 0.9763
Ours ($v=2^{16}$, Exact) 0.9487 0.9436 0.9164 0.8595 0.8706 0.7824 0.4435 0.2688 0.9919 0.9900 0.9825 0.9773
Ours ($v=2^{12}$, KNN) 0.9487 0.9385 0.9062 0.8544 0.9294 0.8882 0.5624 0.3141 0.9919 0.9913 0.9824 0.9774
Ours ($v=2^{16}$, KNN) 0.9487 0.9436 0.9123 0.8608 0.9412 0.9000 0.5553 0.3129 0.9919 0.9916 0.9844 0.9787
@@
Ours ($v=2^{12}$, Exact) 0.8974 0.8641 0.6790 0.4600 0.8375 0.8187 0.3775 0.2181 0.9740 0.9640 0.9463 0.9344
Ours ($v=2^{16}$, Exact) 0.8923 0.8590 0.6821 0.4682 0.8471 0.7706 0.3553 0.2029 0.9740 0.9650 0.9489 0.9359
Ours ($v=2^{20}$, Exact) 0.8923 0.8590 0.6836 0.4687 0.8471 0.7765 0.3612 0.2047 0.9740 0.9659 0.9499 0.9380
Ours ($v=2^{12}$, KNN) 0.8923 0.8615 0.6805 0.4713 0.8500 0.8125 0.4000 0.2275 0.9745 0.9663 0.9498 0.9413
Ours ($v=2^{16}$, KNN) 0.8923 0.8615 0.6949 0.4836 0.8375 0.8187 0.3775 0.2181 0.9810 0.9671 0.9520 0.9435
[…]
@@
KNN. We utilized hierarchical navigable small world (HNSW) algorithm [27] for KNN in our experiment, and provide the results on other KNN algorithms in Appendix 9.
@@ 4.1 Performance on Conditional Diffusion Models
Results. We report the average detection rate for each test set of subsets in Table 2. Compared to the baselines, our proposed DMin obtains the best performance on all subsets. Compared to the baselines, our proposed DMin achieves the best performance across all subsets. The detection rates for top-50 and top-100 on Lego Sets are lower because the Lego Sets training dataset contains only 40 samples (0.43% of the total). Across all subsets and different values of $k$, the KNN with $v=2^{16}$ achieves the best performance in most cases. Additionally, compared to our method without compression, removing normalization substantially decreases performance, confirming that normalization mitigates the instability of gradients in extremely deep models. Interestingly, KNN search often outperforms exact inner product computation in our experiments across all models and subsets. This improvement is likely due to KNN’s ability to approximate the search process, capturing a broader and more representative subset of neighbors.

### v3

# DMin: Scalable Training Data Influence Estimation for Diffusion Models

@@ Abstract
Identifying the training data samples that most influence a generated image is a critical task in understanding diffusion models (DMs), yet existing influence estimation methods are constrained to small-scale or LoRA-tuned models due to computational limitations. To address this challenge, we propose DMin (Diffusion Model influence), a scalable framework for estimating the influence of each training data sample on a given generated image. To the best of our knowledge, it is the first method capable of influence estimation for DMs with billions of parameters. Leveraging efficient gradient compression, DMin reduces storage requirements from hundreds of TBs to mere MBs or even KBs, and retrieves the top-$k$ most influential training samples in under 1 second, all while maintaining performance. Our empirical results demonstrate DMin is both effective in identifying influential training samples and efficient in terms of computational and storage requirements.
Figure 1: Examples of influential training samples, with prompts displayed below generated image. (SD 3 Medium with LoRA, $v=2^{16}$).
@@ 1 Introduction
Diffusion models have emerged as powerful generative models, capable of producing high-quality images and media across various applications [8, 36, 23, 38]. Despite their impressive performance, the extremely large scale and complexity of the datasets used for training are often sourced broadly from the internet [33, 35, 34]. This vast dataset diversity allows diffusion models to generate an extensive range of content, enhancing their versatility and adaptability across multiple domains [19, 5]. However, it also means that these models may inadvertently generate unexpected or even harmful content, reflecting biases or inaccuracies present in the training data.
This raises an important question: given a generated image, can we estimate the influence of each training data sample on this image? Such an estimation is crucial for various applications, such as understanding potential biases [17, 24] and improving model transparency by tracing the origins of specific generated outputs [16, 7, 13].
Recently, many studies have explored influence estimation in diffusion models [26, 18, 26, 27, 11]. These methods assign an influence score to each training data sample relative to a generated image, quantifying the extent to which each sample impacts the generation process. For instance, DataInf [18] and K-FAC [26] are influence approximation techniques tailored for diffusion models. However, they are both second-order methods that require the inversion of the Hessian matrix. To approximate this inversion, they must load all the gradients of training data samples across several predefined timesteps. Notably, in the case of the full-precision Stable Diffusion 3 medium model [10], the gradient of entire model requires approximately 8 GB of storage. Collecting gradients for one training sample over 10 timesteps would consume $8\times 10=80$ GB. Scaling this requirement to a training dataset of $10,000$ samples results in a storage demand of around 800 TB – far exceeding the capacity of typical memory or even hard drives. Given that diffusion models are often trained on datasets with millions of samples, this storage demand becomes impractical. Consequently, these methods are limited to LoRA-tuned models or small diffusion models [15, 31]. Although some prior works have applied gradient compression, such as SVD [13] and quantization [26], the achieved compression rates are insufficient to maintain performance at this scale.
@@
To address these challenges, in this paper, we propose DMin, a scalable influence estimation framework for diffusion models. Unlike existing approaches that are limited to small models or LoRA-tuned models, the proposed DMin scales effectively to larger diffusion models with billions of parameters. For each data sample, DMin first computes and collects gradients at each timestep, then compresses these gradients to MBs or KBs while maintaining performance. Following this compression, DMin can accurately estimate the influence of each training data sample on a given generated image or retrieve the top-$k$ most influential samples on-the-fly using K-nearest neighbors (KNN) search, enabling further speedup based on the specific task.
@@
  • We introduce DMin, a scalable influence estimation framework for DMs, compatible with various architectures, from small models and LoRA-tuned models to large-scale models with billions of parameters.
@@ 2 Influence Estimation for Diffusion Models
Given a test generation $x^{s}$, where $x^{s}$ is generated by a well-trained diffusion model with parameters $\theta^{*}$, the goal of influence estimation is to estimate the influence of each training data sample $x^{i}(1\leq i\leq N)$ on generating $s$, where $N$ is the size of training dataset. Let $z^{i}_{t}$ represents the latent representation of $x^{i}$ on timestep $t$, $z^{s}_{t}$ denotes the latent representation of test generation $s$ on timestep $t$.
Considering in the $\alpha$-th training iteration, the model parameter $\theta_{\alpha+1}$ are updated from $\theta_{\alpha}$ by gradient descent on the noise prediction loss for batch $B=(B_{z},B_{t},B_{\epsilon})$:
$\theta_{\alpha+1}=\theta_{\alpha}-\eta_{\alpha}\frac{1}{|B|}\sum_{(z_{t},t,\epsilon_{t})\in B}\nabla_{\theta_{\alpha}}\mathcal{L}(f_{\theta_{\alpha}}(z_{t},t),\epsilon_{t})$ (2)
where $\eta_{\alpha}$ denotes the learning rate in the $\alpha$-th iteration, $(z_{t}^{i},t^{i},\epsilon_{t}^{i})\in B$, and the contribution of $(z_{t}^{i},t^{i},\epsilon_{t}^{i})$ to batch gradient is $\frac{1}{|B|}\nabla_{\theta_{\alpha}}\mathcal{L}(f_{\theta_{\alpha}}(z^{i}_{t},t^{i}),\epsilon^{i}_{t})$. The influence of this training iteration $z^{i}_{t}$ with respect to $z^{s}_{t}$ on timestep $t$ can be quantified as the change in loss:
@@
To estimate the influence of training data sample $x^{i}$, we can summing up all the training iteration training on $x^{i}$ and timesteps on $t\in\{1,2,\cdots,T\}$:
@@
where $e$ is the number of epochs, $\bar{\eta}$ is the average learning rate during training, $\epsilon$ corresponds to the Gaussian noise used in the training process. However, storing all Gaussian noise from the training process is impractical. Therefore, we randomize the Gaussian noise following the same distribution as in the training process for influence estimation.
Similarly, for a text-to-image model, the influence of a training data sample $X^{i}=(p^{i},x^{i})$ with respect to test generation $X^{s}=(p^{s},x^{s})$ can be estimated by:
@@ 3.1 Gradient Computation
Since the training dataset remains fixed after training, we can cache the loss gradient of each training data sample, as illustrated in Figure 2(a). For a given training pair $X^{i}=(p^{i},x^{i})$, and a timestep $t$, the training data is processed through the diffusion model in the same way as during training, and a loss is computed between the model-predicted noise and a Gaussian noise $\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$. Back-propagation is then performed to obtain the gradient $g^{i}_{t}$ for the training data pair $X^{i}$ at timestep $t$. Once all gradients $\{g^{i}_{1},g^{i}_{2},\cdots,g^{i}_{T}\}$ for $X^{i}$ at all timesteps are obtained, we apply a compression technique to these gradients and cache the compressed versions for influence estimation. Furthermore, for tasks where only the top-$k$ most influential samples are required, we can construct a KNN index on the compressed gradients to enable efficient querying.
Forward and Backward Passes. In forwarding, following the same process as training, for a training pair $(p^{i},x^{i})$ and a timestep $t$, the prompt $p^{i}$ is passed through the encoder to obtain a prompt embedding, while the image $x^{i}$ is passed through a VAE to obtain a latent representation $z_{0}^{i}$. Gaussian noise is then generated from $\epsilon\sim\mathcal{N}(0,I)$ and added to the latent representation to obtain a noisy latent presentation. Then input the timestep $t$, noisy latent presentation $z^{i}_{t}=z_{0}^{i}+\epsilon$, and the embedding $z_{p}^{i}$ are then fed into the model for the forward pass.
@@
Index Construction for KNN. To further enhance the scalability of DMin, we introduce KNN search for tasks requiring only the top-$k$ most influential samples. After gradient compression, as shown in Figure 2(d), we concatenate all the compressed gradients across timesteps to construct a KNN index, enabling efficient querying during influence estimation. This approach is well-suited for extremely large datasets, allowing for the retrieval of the top-$k$ most influential samples on the fly.
3.2 Influence Estimation
@@ 4 Experiments
Models. For conditional text-to-image diffusion model, we use three different models: (1) SD 1.4 with LoRA [^1] CompVis/stable-diffusion-v1-4, (2) SD 3 Medium with LoRA333stabilityai/stable-diffusion-3-medium-diffusers and (3) SD 3 Medium33footnotemark: 3 (Full parameters). For unconditional diffusion model, we conduct experiments on two Denoising Diffusion Probabilistic Models (DDPM) trained on MNIST2221aurent/ddpm-mnist and CIFAR-10444google/ddpm-cifar10-32. The detailed settings of models are included in Appendix 7.1. We fine-tune models on the combined training dataset mentioned above and evaluate them on the testing dataset. During gradient collection, we collect only the gradients of the parameters in the LoRA components for the LoRA-tuned model, whereas for the fully fine-tuned model, we collect the gradients of all parameters (Figure 2(b)).
@@
Ours (w/o Comp.) 0.8872 0.8359 0.5836 0.3969 0.5647 0.4412 0.1435 0.0894 0.9778 0.9778 0.9911 0.9933
Ours ($v=2^{12}$, Exactly) 0.8667 0.8154 0.5713 0.3836 0.5176 0.3882 0.1435 0.0865 0.9778 0.9889 0.9933 0.9944
Ours ($v=2^{16}$, Exactly) 0.8615 0.8231 0.5718 0.3813 0.5529 0.4353 0.1447 0.0894 0.9778 0.9778 0.9911 0.9933
Ours ($v=2^{20}$, Exactly) 0.8667 0.8154 0.5713 0.3836 0.5647 0.4412 0.1435 0.0894 0.9778 0.9778 0.9911 0.9933
Ours ($v=2^{12}$, KNN) 0.8615 0.8128 0.5405 0.3585 0.5059 0.3647 0.1365 0.0824 0.9778 0.9889 0.9933 0.9944
Ours ($v=2^{16}$, KNN) 0.8615 0.8231 0.5723 0.3808 0.5412 0.4176 0.1388 0.0847 0.9778 0.9889 0.9889 0.9944
@@
Ours (w/o Comp.) 0.9128 0.8974 0.8390 0.7605 0.6118 0.5059 0.2318 0.1488 1.0000 1.0000 1.0000 0.9700
Ours ($v=2^{12}$, Exactly) 0.8974 0.8846 0.8318 0.7608 0.6000 0.5235 0.2306 0.1529 0.9837 0.9835 0.9751 0.9703
Ours ($v=2^{16}$, Exactly) 0.9077 0.8872 0.8405 0.7659 0.5765 0.5118 0.2224 0.1482 0.9848 0.9840 0.9761 0.9718
Ours ($v=2^{20}$, Exactly) 0.9077 0.8872 0.8385 0.7651 0.6000 0.5235 0.2294 0.1506 0.9848 0.9840 0.9762 0.9720
Ours ($v=2^{12}$, KNN) 0.9026 0.8949 0.8415 0.7641 0.7294 0.6529 0.3094 0.1924 0.9854 0.9851 0.9771 0.9717
Ours ($v=2^{16}$, KNN) 0.9128 0.9051 0.8472 0.7721 0.7059 0.6353 0.3035 0.1871 0.9864 0.9862 0.9785 0.9736
@@
Ours ($v=2^{12}$, Exactly) 0.9487 0.9000 0.5385 0.3567 0.5529 0.4412 0.1906 0.1165 0.9882 0.9882 0.9420 0.9063
Ours ($v=2^{16}$, Exactly) 0.9590 0.9308 0.5564 0.3690 0.5765 0.4765 0.2047 0.1282 0.9961 0.9902 0.9514 0.9220
Ours ($v=2^{20}$, Exactly) 0.9641 0.9333 0.5590 0.3708 0.5647 0.4765 0.2071 0.1306 0.9922 0.9922 0.9498 0.9202
Ours ($v=2^{12}$, KNN) 0.9282 0.8641 0.5354 0.3518 0.6125 0.5062 0.2025 0.1288 0.9880 0.9820 0.9472 0.9046
Ours ($v=2^{16}$, KNN) 0.9622 0.9108 0.5622 0.3695 0.6250 0.5437 0.2213 0.1419 0.9960 0.9960 0.9640 0.9308
Table 2: Average detection rates of top-k most influential training data samples. Detection rate=# Samples from Same Subset among Top-k Training $\text{Detection rate}=\frac{\text{\# Samples from Same Subset among Top-k Training Samples}}{k}$ where $k=\{5,10,50,100\}$, indicating the average proportion of samples from the same subset appearing in the top-k influential samples. “Ours (w/o Comp.)” indicates that the gradient vectors are not compressed, while “w/o Norm.” signifies that the gradient vectors are not normalized. “Excatly” denotes exact inner product computation. The results for LiSSA, DataInf and D-TRAK on SD3 Medium (Full) are omitted due to the requirement of over hundreds of TB of cache. Moreover, at this scale, it is impractical for LiSSA and DataInf to approximate the Hessian inversion and for D-TRAK to compute an extremely large random projection matrix.
@@
KNN. We utilized hierarchical navigable small world (HNSW) algorithm [25] for KNN in our experiment, and provide the results of ablation study in Appendix 8.
@@ 4.1 Performance on Conditional Diffusion Models
Results. We report the average detection rate for each test set of subsets in Table 2. Compared to the baselines, our proposed DMin obtains the best performance on all subsets. Compared to the baselines, our proposed DMin achieves the best performance across all subsets. The detection rates for top-50 and top-100 on Lego Sets are lower because the Lego Sets training dataset contains only 40 samples (0.43% of the total). Across all subsets and different values of $k$, $v=2^{16}$ achieves the best performance in most cases, whether using KNN or exact inner product computation. Additionally, compared to our method without compression, removing normalization substantially decreases performance, confirming that normalization mitigates the instability of gradients in extremely deep models. Interestingly, KNN search often outperforms exact inner product computation in our experiments across all models and subsets. This improvement is likely due to KNN’s ability to approximate the search process, capturing a broader and more representative subset of neighbors.